An investment will be worth $1000, $2000, or $5000 at the end of the year. The probabiltiies of the values are 0.25, 0.60, and 0.15 respectively. Determine the mean and variance of the worth of the investment. (You may use Excel to solve this problem).
Answer by Theo(13342)

you could to this manually as shown below.
.25 * 1000 = 250
.60 * 2000 = 1200
.15 * 5000 = 750

sum of multiplied amounts = 2200
sum of multipliers = 1.0
mean is equal to 2200 / 1 = 2200.

sum of squares is equal to:
(1000 - 2200)^2 * .25 = 360,000
(2000 - 2200)^2 * .60 = 24,000
(5000 - 2200)^2 * .15 = 1,176,000
sum = 1,560,000
variance = sum / sum of (.25, .60, .15) = 1) = 1,560,000
standard deviation = sqrt(variance) = 1248.9996.

your solution is that the mean is 2200 and the variance is 1,560,000.

this could be done in excel.
i'm not sure if it can be done directly.
i have the 2007 version which may not allow that that to be done automatically.
i can, however, use excel to create it manually.
this takes the drudgery out of the calculations.

here's what the excel output looks like.